### v1

# Entropy Scaling for Diffusion Coefficients in Mixtures

Abstract
Entropy scaling is a powerful technique that has been used successfully for predicting transport properties of pure components over a wide range of states. However, modeling mixture diffusion coefficients by entropy scaling is an unresolved task. We have tackled this issue and present an entropy scaling framework for predicting mixture self-diffusion as well as mutual diffusion in a thermodynamically consistent way. The predictions of the mixture diffusion coefficients are made based on information on the self-diffusion coefficients of the pure components and the infinite dilution diffusion coefficients. This is accomplished using information on the entropy of the mixture, which is taken here from molecular-based equations of state. The framework is flexible and can be applied together with any equation of state. Examples for the application of the new entropy scaling framework for the prediction of diffusion coefficients in mixtures illustrate its impressive performance. It enables predictions a wide range of temperatures and pressures including gaseous, liquid, supercritical, and metastable states - also for strongly non-ideal mixtures.
I Introduction
Diffusion in mixtures is important in many natural and technical processes. And, numbers for diffusion coefficients are needed in many applications, e.g. for the design of fluid separation processes and reactors or for modeling combustion processes. However, experimental data on diffusion coefficients are notoriously scarce, so that reliable models for their prediction are needed. We introduce a new model type that is based on entropy scaling and enables predictions that were previously infeasible.
In diffusion, two phenomena are distinguished: Self-diffusion (a.k.a. tracer diffusion) and mutual (transport) diffusion. Self-diffusion and the corresponding self-diffusion coefficient $D_{i}$ of a component $i$ describes the Brownian movement of individual particles and is defined for pure components and mixtures. In contrast, mutual diffusion is only defined in mixtures and describes the motion of particle collectives of the different components that results in macroscopic mass transfer. Obviously, self-diffusion and mutual diffusion are closely related, but there exist no generally applicable equations that relate self-diffusion coefficients and mutual diffusion coefficients. The different diffusion coefficients are schematically depicted in Fig. 1 for a binary mixture.
Figure 1: Schematic representation of the diffusion coefficients in a binary mixture as a function of the mole fraction. Lines indicate the self-diffusion coefficient of component $i$ $D_{i}$ (blue) and component $j$ $D_{j}$ (green), the Maxwell-Stefan diffusion coefficient Ð$\textit{\DH}_{ij}$ (dark grey), and the Fickian diffusion coefficient $D_{ij}$ (light grey). The limiting cases $D_{i}^{\rm pure}$, $D_{j}^{\rm pure}$ (both pure cases) and $D_{i}^{\infty}$, $D_{j}^{\infty}$ (both pseudo-pure cases) are also depicted.
For describing mutual diffusion [1], there are two common frameworks: Those of Fick and Maxwell-Stefan. The corresponding diffusion coefficients are the Fickian diffusion coefficients $D_{ij}$ (related to the concentration gradient as driving force) and the Maxwell-Stefan diffusion coefficients Ð$\textit{\DH}_{ij}$ (related to the chemical potential gradient as driving force). The two frameworks are thermodynamically consistent representations and can be transformed into each other. For a binary mixture, these diffusion coefficients are related by
$D_{ij}=\textit{\DH}_{ij}\Gamma_{ij},$ (1)
where $\Gamma_{ij}$ is the thermodynamic factor, which is defined by
$\Gamma_{ij}=\frac{x_{i}x_{j}}{RT}\frac{\partial^{2}G}{\partial x_{i}^{2}},$ (2)
where $x_{i}$ and $x_{j}$ are the mole fractions of components $i$ and $j$, respectively, $R$ is the universal gas constant, and $G$ is the Gibbs energy of the mixture.
Evidently, the two coefficients $D_{ij}$ and Ð$\textit{\DH}_{ij}$ become equal if the thermodynamic factor is unity, which is the case in the infinite dilution limit (and for ideal mixtures). Furthermore, self-diffusion coefficient and the mutual diffusion coefficients are related in the infinite dilution limit (cf. Fig. 1). Thus, the following relations apply for the infinite dilution limit:
$x_{i}\rightarrow 0:\quad$$\textit{\DH}_{ij}=D_{ij}=D_{i}=D_{i}^{\infty}$ (3)
$x_{j}\rightarrow 0:\quad$$\textit{\DH}_{ij}=D_{ij}=D_{j}=D_{j}^{\infty}$ (4)
where $D_{i}^{\infty}$ is the diffusion coefficient of component $i$ infinitely diluted in component $j$, $D_{j}^{\infty}$ the diffusion coefficient of component $j$ infinitely diluted in component $i$. Modeling the different diffusion coefficients in a mixture in a consistent way is a challenging task. In this work, we propose a new methodology that provides such a framework.
Physical models for predicting mixture diffusion coefficients at gaseous states are known and established for a long time within kinetic gas theory [2] - see Supplementary Information for details. The prediction of mixture diffusion coefficients at states where significant intermolecular interactions prevail, on the other hand, is still an unresolved problem. Einstein has proposed a method for estimating infinite dilution diffusion coefficients in liquids [3], for which several modifications exist today [2, 4, 5]. For estimating the concentration dependence of mutual diffusion coefficients in mixtures, several empirical models have been proposed, e.g. the Vignes model and the generalized Darken model [6, 7, 8]. However, these models often fail for strongly non-ideal mixtures (see Supplementary Information).
In recent years, entropy scaling has received significant attention for modeling transport properties [9, 10, 11]. It is based on the discovery of Rosenfeld [12, 13], that dynamic properties (i.e. viscosity, thermal conductivity, and self-diffusion coefficient) of pure components, when properly scaled by the density and temperature, are a monovariate function of the configurational entropy $s_{\rm conf}$. This scaling behavior is physically based and related to isomorph theory [14, 11] (however, it is not yet fully understood). For describing the entropy with respect to the desired state point (e.g. given by $T,p$), usually an equation of state (EOS) is used. Entropy scaling can be favorably combined with molecular-based EOS[15], which enables predictions beyond the available data [16, 17]. Entropy scaling is well established for predicting the viscosity and thermal conductivity of mixtures [18, 19]. The corresponding models are based on combination and mixing rules and often enable a reliable prediction of mixture viscosities and thermal conductivities. However, entropy scaling for mixture diffusion coefficients – as depicted in Fig. 1 – has not yet been developed. So far, only the pure component limiting cases for the self-diffusion coefficients $D_{1}^{\rm pure}$ and $D_{2}^{\rm pure}$ (cf. Fig. 1) can be described by entropy scaling models from the literature[20, 21, 22, 23, 16]. Molecular dynamics simulation results indicate that also mixture diffusion coefficients might be subject to a scaling with respect to the entropy [24, 10, 25]. However, no modeling framework for describing mixture diffusion coefficients has been developed yet.
In this work, we propose an entropy scaling model for predicting mixture diffusion coefficients, namely the self-diffusion coefficients as well as the (Fickian and MS) mutual diffusion coefficients, without any adjustable mixture parameters. The physical framework we have developed for predicting diffusion coefficients in mixtures is unique and can be characterized as follows: (1) It can be applied in the entire fluid region, i.e. it covers gases, liquids, and supercritical fluids, phase equilibria, and even metastable states. (2) It describes both self-diffusion and mutual diffusion in a consistent way. (3) It comprises the dependence of the diffusion coefficients on temperature, pressure and composition in the entire fluid region. The applicability of the model is demonstrated for binary mixtures in this work.
The approach developed in this work is based on three central ideas and concepts: (1) Infinite dilution diffusion coefficients are treated as pseudo-pure components that exhibit a monovariate scaling behavior, which can be treated by classical entropy scaling [16]. (2) Therefore, the mixture diffusion coefficient limiting cases, i.e. the pure component and pseudo-pure component self-diffusion coefficients (cf. Fig. 1), are modeled as functions of the entropy. This requires at least one reference data point for each limiting case. (3) Based on the information of the limiting cases, the concentration dependence of the diffusion coefficients $D_{i}$, $D_{j}$, Ð$\textit{\DH}_{ij}$, and $D_{ij}$ are predicted using combination and mixing rules. We demonstrate that these predictions can succeed without any adjustable mixture parameter. The performance of the new approach is demonstrated using model fluids as well as real substance systems. In this work, we use molecular-based EOS models. Yet, also other EOS types such as multiparameter [26] or cubic EOS [27] could be used.
II Results
First, we demonstrate the applicability of a monovariate scaling to the infinite dilution self-diffusion coefficients by treating $D_{i}^{\infty}$ as a pseudo-pure component. This enables the prediction of $D_{i}^{\infty}$ to practically all fluid states based on very little data. Then, we present results for the prediction of the different mixture diffusion coefficients for selected examples (more given in the Supplementary Information).
II.1 Infinite Dilution Diffusion Coefficients
Fig. 2a demonstrates the monovariate scaling behavior (cf. Eqs. (5) - (7)) for infinite dilution diffusion coefficients $\widehat{D}_{2}^{\infty,\circ}$ for a binary Lennard-Jones mixture.
Figure 2: Entropy scaling of diffusion coefficients in a Lennard-Jones mixture with $\sigma_{2}=\sigma_{1}$, $\varepsilon_{2}=0.9\,\varepsilon_{1}$, and $\varepsilon_{12}=1.2\,\sqrt{\varepsilon_{1}\varepsilon_{2}}$. a) Scaled infinite dilution diffusion coefficient of component 2 $\widehat{D}^{\infty,\circ}_{2}$ as a function the reduced configurational entropy $s_{\rm conf}/R$. The line indicates the entropy scaling model. Symbols are MD simulation data from this work. The inset shows the simulation state points in the temperature-density phase diagram of component 1. Therein, solid lines indicate the phase envelopes from Refs. 28, 29. b) Self-diffusion coefficient of component 2 $D_{2}$ (top) and infinite dilution diffusion coefficient of component 2 $D^{\infty}_{2}$ (bottom) as function of the pressure. Lines are the entropy scaling model and symbols are simulation results from Ref. 16 ($D_{2}^{\rm pure}$) and from this work ($D^{\infty}_{2}$). The black solid lines indicates the vapor-liquid equilibrium and the black dashed line the corresponding spinodal. Colors indicate the temperature.
This is impressive considering the fact that a large range of states was studied, cf. Fig. 2a-inset. The quality of the scaling of $\widehat{D}_{2}^{\infty,\circ}$ (i.e. how well the data can be described by a monovariate function) is essentially the same as that found for the pure component diffusion coefficient $D_{2}$. This supports the picture introduced in this work that $D_{2}^{\infty}$ can be considered as a pseudo-pure component property. This picture is physically meaningful considering the fact that in both cases, the mobility of a single particle in a homogeneous environment is described. The Chapman-Enskog diffusion coefficient agrees well with the low-density simulation results – as indicated by the convergence of $\widehat{D}^{\infty,\circ}_{2}\rightarrow 1$ for $s_{\rm conf}/R\rightarrow 0$. In the Supplementary Information, the results for a second binary Lennard-Jones system and a real substance system are presented that support the findings discussed here. For the two Lennard-Jones model systems, the "global" parameters from Ref. 16 were sufficient for describing $\widehat{D}_{2}^{\infty,\circ}$. This finding was not expected and demonstrates the broad applicability of the universal parameters $g_{1}$ and $g_{2}$ in the correlation, that were determined in Ref. 16.
From the models for the pure component self-diffusion coefficient $\widehat{D}_{2}^{\rm pure}(\tilde{s}_{\rm conf})$ [16] and the infinite dilution diffusion coefficient $\widehat{D}_{2}^{\infty}(\tilde{s}_{\rm conf})$, both $D_{2}^{\rm pure}$ and $D_{2}^{\infty}$ can be predicted in a wide range of states. For the Lennard-Jones mixture, the configurational entropy was taken from the Kolafa-Nezbeda EOS[30]. Fig. 2b shows the results of the entropy scaling models for $D_{2}$ and $D_{2}^{\infty}$ in comparison to the MD results. The agreement between both methods is excellent, which is a result of the good performance of the EOS as well as of the fact that both properties show a monovariate relation with respect to the entropy.
Hence, reliable analytical models for the limiting cases $D_{1}$, $D_{2}$, $D_{1}^{\infty}$, and $D_{2}^{\infty}$ (cf. Fig. 1) are now available that can be applied in all fluid regions. They are the basis for the next step, the prediction of $D_{1}$, $D_{2}$, Ð$\textit{\DH}_{12}$, and $D_{12}$ at arbitrary compositions in the mixture.
II.2 Diffusion Coefficients in Mixtures
The model proposed in this work (see methods section) is able to predict all diffusion coefficients in a mixture, which is demonstrated here for binary systems. Fig. 3 shows the model predictions for all four diffusion coefficients ($D_{1}$, $D_{2}$, Ð$\textit{\DH}_{12}$, $D_{12}$) of the Lennard-Jones mixture for the entire composition range and different temperatures. For comparison, simulation data from Ref. [31] are used. The predictions from the entropy scaling framework and the computer experiment results agree well for all four diffusion coefficients (see Fig. 3).
For both self-diffusion coefficients, for which the simulation results have a smaller statistical uncertainty than for the mutual diffusion coefficients, the entropy scaling model describes most simulation data within their uncertainty, despite the strong non-ideality that leads to extrema at about $x_{2}\approx 0.4\,{\rm mol\,mol^{-1}}$. The simulation data for Ð$\textit{\DH}_{12}$ scatter more and have larger error bars than those for $D_{1}$ and $D_{2}$. The entropy scaling framework predicts the Ð$\textit{\DH}_{12}$ simulation data very well in the entire composition range. In particular, the strongly non-ideal behavior is captured accurately by the entropy scaling framework in contrast to established empirical models such as the Vignes or Darken model that fail here (see Supplementary Information). Also the temperature and pressure dependency of the diffusion coefficients are very well described by the model (cf. Fig. 3 and Supplementary Information, respectively). For the predictions of the Fickian diffusion coefficient, the trends are correctly predicted by the entropy scaling predictions, but deviations are observed, cf. Fig. 3d. The deviations between simulation results and entropy scaling are larger for the Fickian diffusion coefficient, which can be attributed to the inclusion of the thermodynamic factor adding additional complexity to the model predictions (and the MD sampling [25]). In contrast to the other three diffusion coefficients ($D_{1}$, $D_{2}$, Ð$\textit{\DH}_{12}$), the Fickian diffusion coefficient $D_{12}$ exhibits a maximum at about $x_{2}\approx 0.55\,{\rm mol\,mol^{-1}}$, which is in line with values of $\Gamma_{12}\gg 1$ obtained in this system. Due to the coupling of entropy scaling and EOS, the model is inherently consistent and can be applied over a wide range of states. As an example, the results for a Lennard-Jones mixture with $\xi_{12}<1$ that exhibits a liquid-liquid miscibility gap are shown in the Supplementary Information.
Figure 3: Diffusion coefficients in a binary Lennard-Jones mixture with $\sigma_{2}=\sigma_{1}$, $\varepsilon_{2}=0.9\,\varepsilon_{1}$, and $\varepsilon_{12}=1.2\,\sqrt{\varepsilon_{1}\varepsilon_{2}}$ as a function of the mole fraction $x_{2}$ at $p=0.13\,\sigma^{3}_{1}\varepsilon_{1}^{-1}$. a) Self-diffusion coefficient of component 1 $D_{1}$; b) Self-diffusion coefficient of component 2 $D_{2}$; c) Maxwell-Stefan diffusion coefficient Ð$\textit{\DH}_{12}$; d) Fickian diffusion coefficient $D_{12}$. Lines are predictions from the entropy scaling framework. Symbols are simulation results from Ref. 31. The entropy scaling framework was used in combination with the Kolafa-Nezbeda EOS[30]. The colors indicate the temperature $T\in\{0.79,0.855,0.92,0.985,1.05\}\,k_{\rm B}\varepsilon_{1}^{-1}$ (from blue to yellow).
Fig. 4 shows the results for the diffusion coefficients in two real substance systems: A mixture of alkanes, $n$-hexane (C6H14) + $n$-dodecane (C12H26), and the mixture toluene (C7H8) + $n$-hexane (C6H14). For both systems, only little experimental data are available (which is common for practically all real mixtures), which is used for the validation of the predictions. The thermodynamic properties of the systems were modeled by the PC-SAFT EOS [32] and the component-specific models for the pure substances were taken from the literature (see Supplementary Information for details).
Figure 4: Diffusion coefficients of real substance mixtures predicted by entropy scaling. Symbols are experimental data and lines are predictions from the entropy scaling framework. a) Self-diffusion coefficients $D_{1}$ in the mixture $n$-hexane (1) + $n$-dodecane (2) as a function of the mole fraction of $n$-dodecane $x_{2}$ at $p=0.1\,{\rm MPa}$. Experimental data are from Ref. 33. b) Fickian diffusion coefficient $D_{\rm 12}$ in the mixture toluene (1) + $n$-hexane (2) as a function of the mole fraction of $n$-hexane $x_{2}$ at $p=0.1\,{\rm MPa}$. Experimental data are from Ref. 34.
Fig. 4a shows the entropy scaling predictions in comparison to the experimental reference data for both self-diffusion coefficients of the system $n$-hexane + $n$-dodecane at two temperatures. Both components strongly differ in their molar mass. Both self-diffusion coefficients decrease with increasing mole fraction $x_{2}$ in a non-linear way, which is predicted well by the framework. The entropy scaling predictions of both self-diffusion coefficients and the experimental data are in excellent agreement.
Fig. 4b shows the results for the Fickian diffusion coefficient in the system toluene + $n$-hexane $D_{12}$ at four temperatures from 293.15 K to 308.15 K as a function of the mole fraction at ambient pressure. The Fickian diffusion coefficient increases with increasing temperature. Again, the entropy scaling predictions are in excellent agreement with the experimental data.
Fig. 5 shows the predictions for the Maxwell-Stefan diffusion coefficient phase diagrams of the system toluene + $n$-hexane at two pressures: $p=0.1\,{\rm MPa}$ (top) and $p=4\,{\rm MPa}$ (bottom). The diffusion coefficients of the coexisting phases can be obtained in a straightforward manner by the framework. Additional results (including supercritical, metastable, and unstable states) are presented in the Supplementary Information. At $p=0.1\,{\rm MPa}$, both components are subcritical (see inset in Fig. 5 top). The Maxwell-Stefan diffusion coefficient is shown for three isotherms (340 K, 360 K, 385 K). As shown in the inset (see Fig. 5 top), one isotherm is entirely in the liquid phase (340 K), one isotherm passes through the vapor-liquid equilibrium (360 K), and one isotherm is entirely in the gas phase (385 K). The Maxwell-Stefan diffusion coefficient in the gas phase is significantly larger than that in the liquid phase. In the liquid phase, the diffusion coefficient strongly depends on the composition. In the gas phase, the Maxwell-Stefan diffusion coefficient exhibits only a minor dependency on the composition. This is in line with the Chapman-Enskog theory[2] – which is inherently incorporated in our framework. At $p=4\,{\rm MPa}$ (cf. Fig. 5 bottom), $n$-hexane is supercritical and a critical point exists at about $T=547.18\,{\rm K}$ (see inset). Four isotherms are plotted: One subcritical isotherm ($T=540\,{\rm K}$), the critical isotherm ($T=547.18\,{\rm}K$), one isotherm passing through the vapor-liquid equilibrium ($T=570\,{\rm K}$), and one isotherm entirely in the gas/supercritical phase ($T=590\,{\rm K}$). The isotherm at $T=540\,{\rm K}$ exhibits a transition between the two pseudo-pure component limiting cases – from a liquid state at $x_{2}=0$ to a supercritical state at $x_{2}=1\;{\rm mol\,mol^{-1}}$. At the critical point, the diffusion coefficient exhibits a large gradient with respect to the composition. Due to the thermodynamic consistency established by the EOS, the framework correctly predicts the Fickian diffusion coefficient to be zero at critical points (see Supplementary Information).
Figure 5: Maxwell-Stefan diffusion coefficient Ð$\textit{\DH}_{12}$ in the mixture toluene (1) + $n$-hexane (2) as a function of the mole fraction of $n$-hexane $x_{2}$ at $p=0.1\,{\rm MPa}$ (top) and $p=4\,{\rm MPa}$ (bottom). Black lines: Vapor-liquid equilibrium; colored lines: Isotherms. Insets show the $T-x$ vapor liquid-equilibrium phase diagram at the respective pressure.
In the Supplementary Information, it is shown that also diffusion coefficients of metastable and unstable states as well as at spinodal states can be described by the framework (which is for example relevant for nucleation). This significantly exceeds the capabilities of all presently available diffusion coefficient models.
III Discussion
In this work, a new methodology for predicting diffusion coefficients in mixtures in a consistent manner for all fluid states is proposed. This includes self-diffusion and mutual diffusion coefficients. The new approach combines several physical concepts, including the Rosenfeld scaling, the Chapman-Enskog theory, and a molecular-based EOS. The framework enables predictions for the mixture diffusion coefficients without any adjustable parameters based on the limiting cases of the pure components and pseudo-pure components (i.e. infinite dilution). Due to the coupling of entropy scaling and EOS, the framework can consistently describe diffusion coefficients in mixtures in different phases (gas, liquid, supercritical, metastable) – including coexisting phases such as vapor-liquid and liquid-liquid equilibria. Also predictions to regions in which no data are available are possible. The strong predictive capabilities are a result of the physical backbone of the framework. The fact that infinite dilution diffusion coefficients exhibit a monovariate scaling behavior was discovered here for the first time. It enables an efficient entropy scaling modeling of the limiting mixture diffusion coefficient cases based on very few data. Since the limiting case diffusion coefficients are monovariate functions with respect to the entropy, extrapolations beyond the temperature and pressure range where data are available are feasible. In the mixture, no general monovariate behavior retains. Nevertheless, the mixture entropy in combination with mixing and combination rules enables predictions of mixture diffusion coefficients, which was demonstrated here for model mixtures as well as real substance systems.
The new entropy scaling framework requires an accurate and robust EOS model that describes mixture thermodynamic properties reliably, i.e. the phase equilibria, the entropy, the thermodynamic factor, and the second virial coefficient. If such a model is available, the entropy scaling framework proposed in this work is a powerful tool. For future work, the extension to multi-component systems would be interesting.
IV Methods
The framework proposed in this work consists of two new elements: (1) Treating the infinite dilution self-diffusion coefficient as a pseudo pure component to obtain a monovariate scaling; (2) The application of mixing and combination rules for predicting the different diffusion coefficients in a mixture.
IV.1 Entropy Scaling of Infinite Dilution Diffusion Coefficients
The infinite dilution diffusion coefficient $D_{i}^{\infty}(T,p)$ is treated as a pseudo-pure component property such that the corresponding scaled property $\widehat{D}_{i}^{\infty,\circ}(s_{\rm conf})$ exhibits a monovariate function with respect to the configurational entropy $s_{\rm conf}$. For testing this postulate, the ES model for pure component diffusion coefficients from Ref. 16 is adopted here. Thereby, the infinite dilution self-diffusion coefficient is transformed using a modified Rosenfeld scaling (marked by ) as
$D_{i}^{\infty,\circ}=D_{i}^{\infty}\frac{\rho_{\rm N}^{1/3}}{\sqrt{RT/M_{\rm CE}}}\left(\frac{-s_{\rm conf}}{R}\right)^{2/3},$ (5)
where $\rho_{\rm N}$ is the number density of the solvent at given $T$ and $p$. The reference mass $M_{\rm CE}$ is adopted from the Chapman-Enskog (CE) theory as
$M_{\rm CE}=\frac{2}{1/M_{i}+1/M_{j}},$ (6)
where $M_{i}$ and $M_{j}$ are the molar masses of the pure components. For the zero-density limit, the scaled Chapman-Enskog infinite dilution diffusion coefficient $D_{{\rm CE},i}^{\infty,\circ}$ is computed using a Lennard-Jones kernel (see Supplementary Information for details). Therefore, the Lennard-Jones cross-interaction size and energy parameters $\sigma_{ij}$ and $\varepsilon_{ij}$ are calculated by the Lorentz-Berthelot combining rules here [35, 36]. Finally, the CE-scaled infinite dilution diffusion coefficient $\widehat{D}^{\infty,\circ}_{i}$ is described by a split between the low-density (LD) and the high-density (HD) region as
$\widehat{D}^{\infty,\circ}_{i}=\underbrace{\frac{D^{\infty,\circ}_{i}}{D^{\infty,\circ}_{{\rm CE},i}}}_{\rm LD}W(\tilde{s}_{\rm conf})+\underbrace{\frac{D^{\infty,\circ}_{i}}{\min\left(D^{\infty,\circ}_{{\rm CE},i}\right)}}_{\rm HD}(1-W(\tilde{s}_{\rm conf})),$ (7)
where $\tilde{s}_{\rm conf}=-s_{\rm conf}/(Rm)$ is the reduced configurational entropy using the segment parameter $m$ of molecular-based EOS and $W$ is a smoothed step function. The CE-scaled infinite dilution diffusion coefficient $\widehat{D}^{\infty,\circ}_{i}(\tilde{s}_{\rm conf})$ is then described by a function with up to five adjustable parameters $\alpha_{k}$ (see Supplementary Information). In this work, one or two system-dependent parameters were used for the real substance systems. A detailed description of the entropy scaling framework for infinite dilution diffusion coefficients is given in the Supplementary Information. To test and validate this new scaling approach for infinite dilution diffusion coefficients, two model systems and a real substance system (all strongly non-ideal systems) were studied in this work using molecular dynamics simulations using the molecular simulation engine ms2 [37]. Simulations were conducted in the all fluid regions, i.e. liquid, gaseous, supercritical, and metastable states (details are given in the Supplementary Information).
IV.2 Predicting Diffusion Coefficients in Mixtures
For predicting the diffusion coefficients $D_{i}(x_{j})$, $D_{j}(x_{j})$, Ð$\textit{\DH}_{ij}(x_{j})$, and $D_{ij}(x_{j})$ in a mixture, the entropy scaling framework is extended using combining and mixing rules. We briefly outline how this is done here; details are given in the Supplementary Information. Diffusion coefficients in mixtures are predicted based on the limiting case models for the self-diffusion coefficient of pure components and the (infinite dilution) pseudo-pure component. The scaled diffusion coefficient $\widehat{\Lambda}^{\circ}\in\{\widehat{D}_{i}^{\circ},\widehat{D}_{j}^{\circ},\widehat{\textit{\DH}}_{ij}^{\circ}\}$ is computed using the configurational entropy of the mixture $s_{\rm conf}^{\rm mix}=s_{\rm conf}^{\rm mix}(T,\rho,\underline{x})$ and a mixture model for the scaled function $\widehat{\Lambda}^{\circ}=\widehat{\Lambda}^{\circ}(s_{\rm conf}^{\rm mix},\alpha_{k,{\rm mix}})$, which is computed from the limiting case scaling functions. The final (unscaled) diffusion coefficient $\Lambda\in\{D_{i},D_{j},\textit{\DH}_{ij}\}$ is calculated as
$\Lambda=\frac{\widehat{\Lambda}^{\circ}}{\frac{W(\tilde{s}_{\rm conf}^{\rm mix})}{\Lambda^{\circ}_{\rm CE}}+\frac{1-W(\tilde{s}_{\rm conf}^{\rm mix})}{\min\left(\Lambda^{\circ}_{\rm CE}\right)}}\cdot\frac{\sqrt{RT/M_{\rm ref}}}{\rho_{\rm N}^{1/3}}\left(\frac{-s_{\rm conf}^{\rm mix}}{R}\right)^{-2/3},$ (8)
where $\Lambda^{\circ}_{\rm CE}$ is the Chapman-Enskog property of the mixture, $\rho_{\rm N}$ is the number density of the mixture, and $M_{\rm ref}$ is the reference mass of the mixture. The Fickian diffusion coefficient is obtained from the Maxwell-Stefan diffusion coefficient by the thermodynamic factor as predicted by the EOS (see Eq. 1). Details on the calculation of the Chapman-Enskog property of the mixture $\Lambda^{\circ}_{\rm CE}$, the reference mass $M_{\rm ref}$, and $\widehat{\Lambda}^{\circ}=\widehat{\Lambda}^{\circ}(s_{\rm conf}^{\rm mix},\alpha_{k,{\rm mix}})$ are given in the Supplementary Information. All required quantities are obtained using predictive combination rules and mixing rules and the EOS mixture model. No adjustable mixture parameters are introduced. The non-ideality of the mixture is primarily taken into account by the underlying EOS via the predicted mixture entropy. The combination of the EOS and entropy scaling enables the consistent calculation of the diffusion coefficients ($D_{i}$, $D_{j}$, Ð$\textit{\DH}_{ij}$, $D_{ij}$), homogeneous bulk properties ($pvT$, $s_{\rm conf}$, $\Gamma_{ij}$, etc.), and phase equilibria (e.g. vapor-liquid and liquid-liquid equilibria).
Supplementary Information
Supplementary Information with additional results and technical details is available.
Competing Interests
The authors declare no competing interests.
Acknowledgements.
We gratefully acknowledge funding from the European Union’s Horizon Europe research and innovation program under grant agreement no. 101137725 (BatCAT) and from the German Research Foundation (DFG) under grant 548115878. The simulations were carried out on the HPC machine ELWE at the RHRZ under the grant RPTU-MTD and on the HPC machine MOGON at the NHR SW under the grant TU-MSG (supported by the Federal Ministry of Education and Research and the state governments).